OpenStudy (jessie12):

In Mrs. Brown's Geometry class there are 15 boys and 10 girls. The students are presenting projects to the class. If Mrs. Brown selects students at random for presentations, what is the probability that the first 2 students chosen are girls

OpenStudy (anonymous):

\[\frac{10}{25}\times \frac{9}{24}\]